# Towards Implicit Aggregation: Robust Image Representation for Place Recognition in the Transformer Era

### v1

@@ 1 Introduction
Visual place recognition (VPR) involves identifying the coarse geographical location of a query place image by retrieving the most similar images from a geo-tagged database captured at previously visited places survey. It is a fundamental and essential task in a wide range of computer vision and robotics applications, e.g., augmented reality arloc, autonomous driving autodriving, and SLAM slam. Thus, it has garnered significant attention and study. Despite recent advances, there still exist some challenges in VPR, including condition variations, viewpoint changes, and perceptual aliasing (images from different places showing high similarity) survey, etc.
Typically, VPR is formulated as an image retrieval problem. For a given query image and a database, all place images are represented using global features, and the nearest neighbor search is conducted in this feature space to get the target place images that best match the query. The global features are usually obtained by employing aggregation methods (e.g., VLAD vlad2010) to process local features. With the advancement of deep learning, most VPR methods have used a convolutional neural network (CNN) resnet or vision transformer (ViT) vit as the backbone to extract local (patch) features. Meanwhile, NetVLAD netvlad and GeM pooling gem have become the most popular aggregation methods for aggregating local features to yield global descriptors, which are generally robust against common visual variations. Following this paradigm, some recent studies proposed more aggregation methods (e.g., MixVPR mixvpr, SALAD salad, CricaVPR cricavpr, BoQ boq, and EDTformer edtformer), trying to make the global features condition- and viewpoint-invariant, thereby achieving a promising performance.
@@
Although this backbone-plus-aggregator VPR paradigm to obtain global features has become the de-facto standard benchmark in the CNN era, it has some potential issues. First, the two-stage process (feature extraction + aggregation) may lead to unnecessary structural complexity and redundancy. Second, the one-shot aggregation of patch features implemented by the aggregator offers no opportunity for correction and refinement. Regarding specific aggregation methods (aggregators), there may exist some particular issues, such as the loss of position information of original patch features in NetVLAD netvlad. Designing a perfect aggregator artificially is highly challenging. However, in light of the nature of transformer-based backbones, which are capable of modeling global contextual information and long-range dependencies yolos, we argue that it is no longer necessary to design an aggregator separately. Instead, we can leverage the intrinsic self-attention mechanism within the backbone to implicitly aggregate useful information from patch tokens, thereby eliminating the need for an extra aggregator. In fact, previous work BoQ boq also attempted to utilize self- and cross-attention mechanisms to aggregate useful information, yet it still introduced an extra aggregator that includes encoder blocks and cross-attention layers, as well as a large number of learnable queries. Another study dinoregisters indicated that simply adding some registers, similar to concatenating the class token with patch tokens, can buffer excess global information (so-called "undesirable artifacts") into these registers. Unfortunately, it discarded these registers finally and lacked deeper research on the use of global information on registers.
@@ 2 Related Work
Visual Place Recognition: Early research on VPR primarily focused on aggregating the hand-crafted descriptors SURF to global descriptors using some classical aggregation algorithms, such as Bags of Words BoW and VLAD vlad2010; VLAD1; densevlad; allVLAD; pbVLAD. In light of the remarkable achievements of deep learning across numerous computer vision tasks, contemporary VPR approaches sunderhaufIROS2015; netvlad; crn; landmarks2; semantic; categorization1; categorization2; landmarks3; yin2019; sfrs; eigenplaces; nprliu have increasingly utilized diverse deep features for better performance. Besides, traditional aggregation algorithms are gradually replaced by trainable aggregation layers, e.g., NetVLAD netvlad and GeM pooling gem. Although some methods patchvlad; geowarp; shen2023structvpr; dhevpr; selavpr employ local feature matching for re-ranking after initial global feature retrieval to boost performance, the backbone-plus-aggregator paradigm has been the de-facto standard benchmark in VPR over the past decade. Some recent research cosplace; gsv; mixvpr; salad; boq; cricavpr has proposed several alternative approaches following this paradigm. For instance, CricaVPR cricavpr leveraged a cross-image encoder to produce cross-image correlation-aware global representations. SALAD salad redefined the soft assignment in NetVLAD as an optimal transport problem and used the Sinkhorn algorithm to solve it. BoQ boq employed distinct learnable queries to probe the input features through cross-attention, facilitating better information aggregation. These methods achieved excellent results using the ViT-based foundation model DINOv2 as the backbone. Unlike these methods that meticulously design auxiliary aggregators to yield global features, our ImAge method presents a novel paradigm that only introduces some additional tokens to the transformer backbone to conduct implicit aggregation via the inherent self-attention mechanism in transformers, thus achieving a simpler architecture and more powerful performance.
@@
Additional Tokens in Transformers: Popularized by BERT (devlin2019bert), integrating special tokens into the token sequence in transformers has been a promising design choice for various purposes. We group such extra tokens into 3 categories based on their functional roles. 1) Output-oriented tokens are learnable anchors that collect information from patch tokens, whose output values are then transmitted as task-specific outputs, e.g., the class tokens used in BERT devlin2019bert and ViT vit for classification, as well as detection tokens in YOLOS fang2021you for object detection. 2) Prompt tokens act as trainable continuous vectors that replace traditional discrete text prompts, efficiently guiding pretrained transformer models to adapt to specific tasks by adjusting the model input, without modifying the parameters of models liPrefixTuningOptimizingContinuous2021; prompttuning; jiaVisualPromptTuning2022, which has become an essential branch of parameter-efficient fine-tuning methods houlsbyParameterEfficientTransferLearning. 3) Memory tokens act as registers that hold intermediate states during sequential processing steps, tracing their roots to neural memory architectures burtsev2020memory; bulatov2022recurrent. This approach gains critical support from the DINOv2-register work dinoregisters, which observed that vision transformers improperly re-purpose background patch tokens as implicit memories when the standard class token lacks the capacity to accommodate global semantics. To address this, they prepend multiple memory tokens called registers to input tokens, which provide extra storage for buffering of global context, thus eliminating artifacts. Inspired by this work, we introduce the concept of aggregation tokens to effectively absorb global context from patch tokens. However, register tokens are discarded from the final output after temporary use, contrasting with our aggregation tokens that directly form the output descriptor for VPR (i.e., our method falls into the "output-oriented tokens" category). Among VPR methods, BoQ boq also advocated the introduction of a bag of output-oriented tokens named queries for aggregation (but in the aggregator rather than backbone). While effective, BoQ uses extra encoder blocks and cross-attention layers as the aggregator. In contrast, our method directly employs the inherent self-attention mechanism of the backbone, offering unique advantages.
@@ 3 Methodology
This section begins with a review of the ViT vit and the self-attention mechanism in it, which serves as the foundation for our ImAge method. Following that, we first present the pipeline of our method. Then, we introduce the insertion strategy of our aggregation tokens and their initialization method.
@@ 3.1 Preliminary
ViT vit and its variants have rapidly emerged as the preferred backbones for a variety of computer vision tasks yue2025learning; zhangquan2024distilling; 3dpose2024; 3dpose2025; selavpr, owing to their exceptional capacity for modeling global relationships raghu2021vision. Given an input image of size $H\times W$, ViT partitions it into $N={HW}/{P^{2}}$ non-overlapping patches. Each patch is then flattened and linearly projected to create a $D$-dimensional token $x_{p}^{i}$. A learnable class token $x_{\mathsf{CLS}}\in\mathbb{R}^{D}$ is prepended to this sequence, and positional embeddings are added to encode spatial information, forming the initial input token sequence $z_{0}=[x_{\mathsf{CLS}},x_{p}^{1},\dots,x_{p}^{N}]\in\mathbb{R}^{(N+1)\times D}$. This sequence is iteratively processed through $L$ transformer encoder blocks. Each block comprises three core components: layer normalization (LN), multi-head self-attention (MHSA), and multi-layer perceptron (MLP). The $l$-th block updates the input $z_{l-1}$ to $z_{l}$ via
@@ 3.2 Implicit Aggregation via the Transformer Backbone
After extracting the patch features/tokens via the backbone, there are primarily two ways in previous works to obtain robust global descriptors. One is to directly aggregate these patch tokens with a common aggregator (e.g., NetVLAD netvlad and SALAD salad), as in Fig. 2 (a). The other uses the queries-based aggregator to learn global information from the patch tokens (e.g., BoQ boq and EDTformer edtformer), as in Fig. 2 (b). However, our ImAge will essentially eliminate the use of aggregators.
@@
Obviously, our ImAge fundamentally diverges from the practices of prompt tuning (aim to fine-tune models) prompttuning and register tokens (aim to remove artifacts) dinoregisters, which discard the newly added tokens finally. Besides, our method also differs from the class token. Our agg tokens have better scalability, along with different insertion strategies and initialization methods, which will be described below.
@@ 3.3 The Insertion Strategy of Aggregation Tokens
Our implicit aggregation method provides a robust image representation for VPR in an extremely simple manner. It requires neither explicit aggregators nor any modifications to the original backbone. However, where and how to add our agg tokens remains an open issue worthy of further exploration. For instance, previous works such as prompt tuning and DINOv2-register prepend additional tokens to the patch tokens (and class token) before the first transformer block, as shown in Fig. 3 (a). Our objective differs from these works, and we no longer follow this way regarding the specific placement of agg tokens. More specifically, there are two reasons for this: 1) Our goal is to aggregate patch tokens with meaningful representations. Since early transformer blocks produce relatively weak features, adding agg tokens at the beginning is unnecessary and potentially detrimental to model performance. 2) In the field of VPR, the common practice for model training is to fine-tune only the last few blocks (layers) of the pre-trained model on the VPR dataset, while freezing the previous blocks. If agg tokens are added at the beginning, although most of the shallow and middle blocks are frozen, the added agg tokens need to be tuned. According to the chain rule of back-propagation losa, the gradients of the parameters in these frozen blocks still need to be calculated, leading to significant GPU memory and computational overhead.
In light of the above considerations, our strategy is to prepend agg tokens only when the patch tokens have acquired sufficient representational capability. A more specific criterion is to add the agg tokens at the junction between frozen and trainable transformer blocks, as illustrated in Fig. 3 (b). For example, in the case of the DINOv2 backbone, most previous VPR methods salad; SuperVLAD only fine-tune the last four blocks. Accordingly, we prepend the agg tokens to the patch tokens before the fourth-to-last block. Since the preceding blocks are frozen, it indicates that the features output here are general enough. The subsequent blocks are then trained on the VPR dataset to produce features more suitable for the VPR task, so our agg tokens can also learn better task-specific global representations. Additionally, we consider two alternative strategies. One is to add agg tokens before a deeper trainable block, as shown in Fig. 3 (c). The other is to add agg tokens progressively instead of all at once, as shown in Fig. 3 (d). However, both ways reduce the opportunities for the refinement and correction of image representations, which leads to suboptimal performance. Based on these objective factors, we finally propose the aforementioned strategy (b) for the insertion of agg tokens.
@@ 3.4 The Initialization of Aggregation Tokens
Fortunately, a similar issue has been discussed in NetVLAD NetVLADpami. This method determines $k$ cluster centers (and the parameters of the assignment layer) through training. The residual statistics from patch features to cluster centers are used as the global representation. At the beginning, it also requires initializing $k$ cluster centers and the soft-assignment layer through the unsupervised $k$-means algorithm to achieve good performance. Although our ImAge method uses the self-attention mechanism to perform implicit aggregation, its essence can be regarded as each added agg token representing a unique category (but not necessarily corresponding to an object category in human semantics, such as building or vegetation) that is helpful to VPR, similar to each cluster in NetVLAD. Therefore, we can learn from NetVLAD, using the $k$-means algorithm to perform unsupervised clustering for the initialization of agg tokens. Besides, NetVLAD uses L2-normalized cluster centers to initialize parameters (weight $\mathbf{w}$) in the assignment layer. Through our empirical research, the L2-normalized centers can reduce the impact of extreme cases and are more suitable for initializing agg tokens than the original centers, i.e., it is our final method.
@@ 4.1 Datasets and Performance Evaluation
Datasets. We conduct the experiments on several VPR benchmark datasets. These datasets exhibit various challenges, including viewpoint changes, condition variations, and the perceptual aliasing issue. Table 1 provides a summary of the main evaluation datasets. MSLS msls is a particularly challenging dataset, in which images are taken from urban, suburban, and natural scenes, covering diverse visual changes. Pitts30k pitts extracted from Google Street View, mainly presents severe variations in viewpoint. Tokyo24/7 densevlad shows dramatic condition (light) changes. Nordland nordland is gathered across four seasons with a fixed perspective from the front of a train. Moreover, we also use the Baidu Mall baidumall, SPED SPED, Pitts250k pitts, St. Lucia stlucia, and SVOX SVOX datasets for a few supplementary experiments.
@@
Performance Evaluation. We follow the previous work cosplace; benchmark using the Recall@N (R@N) as the evaluation metric for recognition performance. R@N is the proportion of queries for which at least one of the top-N predicted images is within a threshold of ground truth. We set the threshold to 10 frames for Nordland and 25 meters for others, as in this benchmark benchmark.

### v2

@@ 1 Introduction
Visual place recognition (VPR) involves identifying the coarse geographical location of a query place image by retrieving the most similar images from a geo-tagged database captured at previously visited places Lowry et al. (2015). It is a fundamental and essential task in a wide range of computer vision and robotics applications, e.g., augmented reality Middelberg et al. (2014), autonomous driving Doan et al. (2019), and SLAM Campos et al. (2021). Thus, it has garnered significant attention and study. Despite recent advances, there still exist some challenges in VPR, including condition variations, viewpoint changes, and perceptual aliasing (images from different places showing high similarity) Lowry et al. (2015), etc.
Typically, VPR is formulated as an image retrieval problem. For a given query image and a database, all place images are represented using global features, and the nearest neighbor search is conducted in this feature space to get the target place images that best match the query. The global features are usually obtained by employing aggregation methods (e.g., VLAD Jégou et al. (2010)) to process local features. With the advancement of deep learning, most VPR methods have used a convolutional neural network (CNN) He et al. (2016) or vision transformer (ViT) Dosovitskiy et al. (2020) as the backbone to extract local (patch) features. Meanwhile, NetVLAD Arandjelovic et al. (2016) and GeM pooling Radenović et al. (2018) have become the most popular aggregation methods for aggregating local features to yield global descriptors, which are generally robust against common visual variations. Following this paradigm, some recent studies proposed more aggregation methods (e.g., MixVPR Ali-Bey et al. (2023), SALAD Izquierdo and Civera (2024b), CricaVPR Lu et al. (2024b), BoQ Ali-bey et al. (2024), and EDTformer Jin et al. (2025)), trying to make the global features condition- and viewpoint-invariant, thereby achieving a promising performance.
@@
Although this backbone-plus-aggregator VPR paradigm to obtain global features has become the de-facto standard Berton et al. (2022b) in the CNN era, it has some potential issues. First, the two-stage process (feature extraction + aggregation) may lead to unnecessary structural complexity and redundancy. Second, the one-shot aggregation of patch features implemented by the aggregator offers no opportunity for correction and refinement. Regarding specific aggregation methods (aggregators), there may exist some particular issues, such as the loss of position information of original patch features in NetVLAD Arandjelovic et al. (2016). Designing a perfect aggregator artificially is highly challenging. However, in light of the nature of transformer-based backbones, which are capable of modeling global contextual information and long-range dependencies Fang et al. (2021b), we argue that it is no longer necessary to design an aggregator separately. Instead, we can leverage the intrinsic self-attention mechanism within the backbone to implicitly aggregate useful information from patch tokens, thereby eliminating the need for an extra aggregator. In fact, previous work BoQ Ali-bey et al. (2024) also attempted to utilize self- and cross-attention mechanisms to aggregate useful information, yet it still introduced an extra aggregator that includes encoder blocks and cross-attention layers, as well as a large number of learnable queries. Another study Darcet et al. (2024) indicated that simply adding some registers, similar to concatenating the class token with patch tokens, can buffer excess global information (so-called "undesirable artifacts") into these registers. Unfortunately, it discarded these registers finally and lacked deeper research on the use of global information on registers.
@@ 2 Related Work
Visual Place Recognition: Early research on VPR primarily focused on aggregating the hand-crafted descriptors Bay et al. (2008) to global descriptors using some classical aggregation algorithms, such as Bags of Words Sivic and Zisserman (2003) and VLAD Jégou et al. (2010); Lowry and Andreasson (2018); Torii et al. (2015); Arandjelovic and Zisserman (2013); Kim et al. (2015). In light of the remarkable achievements of deep learning across numerous computer vision tasks, contemporary VPR approaches Sünderhauf et al. (2015); Arandjelovic et al. (2016); Jin Kim et al. (2017); Chen et al. (2017b); Naseer et al. (2017); Garg et al. (2017, 2018); Xin et al. (2019); Yin et al. (2019); Ge et al. (2020); Berton et al. (2023); Liu et al. (2024); Dong et al. (2025) have increasingly utilized diverse deep features for better performance. Besides, traditional aggregation algorithms are gradually replaced by trainable aggregation layers, e.g., NetVLAD Arandjelovic et al. (2016) and GeM pooling Radenović et al. (2018). Although some methods Hausler et al. (2021); Berton et al. (2021a); Shen et al. (2023); Lu et al. (2024a, c) employ local feature matching for re-ranking after initial global feature retrieval to boost performance, the backbone-plus-aggregator paradigm has been the de-facto standard Berton et al. (2022b) in VPR over the past decade. Some recent research Berton et al. (2022a); Ali-bey et al. (2022); Ali-Bey et al. (2023); Izquierdo and Civera (2024b); Ali-bey et al. (2024); Lu et al. (2024b) has proposed several alternative approaches following this paradigm. For instance, CricaVPR Lu et al. (2024b) leveraged a cross-image encoder to produce cross-image correlation-aware global representations. SALAD Izquierdo and Civera (2024b) redefined the soft assignment in NetVLAD as an optimal transport problem and used the Sinkhorn algorithm to solve it. BoQ Ali-bey et al. (2024) employed distinct learnable queries to probe the input features through cross-attention, facilitating better information aggregation. These methods achieved excellent results using the ViT-based foundation model DINOv2 as the backbone. Unlike these methods that meticulously design auxiliary aggregators to yield global features, our ImAge method presents a novel paradigm that only introduces some additional tokens to the transformer backbone to conduct implicit aggregation via the inherent self-attention mechanism in transformers, thus achieving a simpler architecture and more powerful performance.
@@
Additional Tokens in Transformers: Popularized by BERT (Devlin et al., 2019), integrating special tokens into the token sequence in transformers has been a promising design choice for various purposes. We group such extra tokens into 3 categories based on their functional roles. 1) Output-oriented tokens are learnable anchors that collect information from patch tokens, whose output values are then transmitted as task-specific outputs, e.g., the class tokens used in BERT Devlin et al. (2019) and ViT Dosovitskiy et al. (2020) for classification, as well as detection tokens in YOLOS Fang et al. (2021a) for object detection. 2) Prompt tokens act as trainable continuous vectors that replace traditional discrete text prompts, efficiently guiding pretrained transformer models to adapt to specific tasks by adjusting the model input, without modifying the parameters of models Li and Liang (2021); Lester et al. (2021); Jia et al. (2022), which has become an essential branch of parameter-efficient fine-tuning methods Houlsby et al. (2019). 3) Memory tokens act as registers that hold intermediate states during sequential processing steps, tracing their roots to neural memory architectures Burtsev et al. (2020); Bulatov et al. (2022). This approach gains critical support from the DINOv2-register work Darcet et al. (2024), which observed that vision transformers improperly re-purpose background patch tokens as implicit memories when the standard class token lacks the capacity to accommodate global semantics. To address this, they prepend multiple memory tokens called registers to input tokens, which provide extra storage for buffering of global context, thus eliminating artifacts. Inspired by this work, we introduce the concept of aggregation tokens to effectively absorb global context from patch tokens. However, register tokens are discarded from the final output after temporary use, contrasting with our aggregation tokens that directly form the output descriptor for VPR (i.e., our method falls into the "output-oriented tokens" category). Among VPR methods, BoQ Ali-bey et al. (2024) also advocated the introduction of a bag of output-oriented tokens named queries for aggregation (but in the aggregator rather than backbone). While effective, BoQ uses extra encoder blocks and cross-attention layers as the aggregator. In contrast, our method directly employs the inherent self-attention mechanism of the backbone, offering unique advantages.
@@ 3 Methodology
This section begins with a review of the ViT Dosovitskiy et al. (2020) and the self-attention mechanism in it, which serves as the foundation for our ImAge method. Following that, we first present the pipeline of our method. Then, we introduce the insertion strategy of our aggregation tokens and their initialization method.
@@ 3.1 Preliminary
ViT Dosovitskiy et al. (2020) and its variants have rapidly emerged as the preferred backbones for a variety of computer vision tasks Yue et al. (2025); Zhang et al. (2024b, a, 2025); Lu et al. (2024c), owing to their exceptional capacity for modeling global relationships Raghu et al. (2021). Given an input image of size $H\times W$, ViT partitions it into $N={HW}/{P^{2}}$ non-overlapping patches. Each patch is then flattened and linearly projected to create a $D$-dimensional token $x_{p}^{i}$. A learnable class token $x_{\mathsf{CLS}}\in\mathbb{R}^{D}$ is prepended to this sequence, and positional embeddings are added to encode spatial information, forming the initial input token sequence $z_{0}=[x_{\mathsf{CLS}},x_{p}^{1},\dots,x_{p}^{N}]\in\mathbb{R}^{(N+1)\times D}$. This sequence is iteratively processed through $L$ transformer encoder blocks. Each block comprises three core components: layer normalization (LN), multi-head self-attention (MHSA), and multi-layer perceptron (MLP). The $l$-th block updates the input $z_{l-1}$ to $z_{l}$ via
@@ 3.2 Implicit Aggregation via the Transformer Backbone
After extracting the patch features/tokens via the backbone, there are primarily two ways in previous works to obtain robust global descriptors. One is to directly aggregate these patch tokens with a common aggregator (e.g., NetVLAD Arandjelovic et al. (2016) and SALAD Izquierdo and Civera (2024b)), as in Fig. 2 (a). The other uses the queries-based aggregator to learn global information from the patch tokens (e.g., BoQ Ali-bey et al. (2024) and EDTformer Jin et al. (2025)), as in Fig. 2 (b). However, our ImAge will essentially eliminate the use of aggregators.
@@
Obviously, our ImAge fundamentally diverges from the practices of prompt tuning (aim to fine-tune models) Lester et al. (2021) and register tokens (aim to remove artifacts) Darcet et al. (2024), which discard the newly added tokens finally. Besides, our method also differs from the class token. Our agg tokens have better scalability, along with different insertion strategies and initialization methods, which will be described below.
@@ 3.3 The Insertion Strategy of Aggregation Tokens
Our implicit aggregation method provides a robust image representation for VPR in an extremely simple manner. It requires neither explicit aggregators nor any modifications to the original backbone. However, where and how to add our agg tokens remains an open issue worthy of further exploration. For instance, previous works such as prompt tuning and DINOv2-register prepend additional tokens to the patch tokens (and class token) before the first transformer block, as shown in Fig. 3 (a). Our objective differs from these works, and we no longer follow this way regarding the specific placement of agg tokens. More specifically, there are two reasons for this: 1) Our goal is to aggregate patch tokens with meaningful representations. Since early transformer blocks produce relatively weak features, adding agg tokens at the beginning is unnecessary and potentially detrimental to model performance. 2) In the field of VPR, the common practice for model training is to fine-tune only the last few blocks (layers) of the pre-trained model on the VPR dataset, while freezing the previous blocks. If agg tokens are added at the beginning, although most of the shallow and middle blocks are frozen, the added agg tokens need to be tuned. According to the chain rule of back-propagation Mercea et al. (2024), the gradients of the parameters in these frozen blocks still need to be calculated, leading to significant GPU memory and computational overhead.
In light of the above considerations, our strategy is to prepend agg tokens only when the patch tokens have acquired sufficient representational capability. A more specific criterion is to add the agg tokens at the junction between frozen and trainable transformer blocks, as illustrated in Fig. 3 (b). For example, in the case of the DINOv2 backbone, most previous VPR methods Izquierdo and Civera (2024b); Lu et al. (2024d) only fine-tune the last four blocks. Accordingly, we prepend the agg tokens to the patch tokens before the fourth-to-last block. Since the preceding blocks are frozen, it indicates that the features output here are general enough. The subsequent blocks are then trained on the VPR dataset to produce features more suitable for the VPR task, so our agg tokens can also learn better task-specific global representations. Additionally, we consider two alternative strategies. One is to add agg tokens before a deeper trainable block, as shown in Fig. 3 (c). The other is to add agg tokens progressively instead of all at once, as shown in Fig. 3 (d). However, both ways reduce the opportunities for the refinement and correction of image representations, which leads to suboptimal performance. Based on these objective factors, we finally propose the aforementioned strategy (b) for the insertion of agg tokens.
@@ 3.4 The Initialization of Aggregation Tokens
Fortunately, a similar issue has been discussed in NetVLAD Arandjelović et al. (2018). This method determines $k$ cluster centers (and the parameters of the assignment layer) through training. The residual statistics from patch features to cluster centers are used as the global representation. At the beginning, it also requires initializing $k$ cluster centers and the soft-assignment layer through the unsupervised $k$-means algorithm to achieve good performance. Although our ImAge method uses the self-attention mechanism to perform implicit aggregation, its essence can be regarded as each added agg token representing a unique category (but not necessarily corresponding to an object category in human semantics, such as building or vegetation) that is helpful to VPR, similar to each cluster in NetVLAD. Therefore, we can learn from NetVLAD, using the $k$-means algorithm to perform unsupervised clustering for the initialization of agg tokens. Besides, NetVLAD uses L2-normalized cluster centers to initialize parameters (weight $\mathbf{w}$) in the assignment layer. Through our empirical research, the L2-normalized centers can reduce the impact of extreme cases and are more suitable for initializing agg tokens than the original centers, i.e., it is our final method.
@@ 4.1 Datasets and Performance Evaluation
Datasets. We conduct the experiments on several VPR benchmark datasets. These datasets exhibit various challenges, including viewpoint changes, condition variations, and the perceptual aliasing issue. Table 1 provides a summary of the main evaluation datasets. MSLS Warburg et al. (2020) is a particularly challenging dataset, in which images are taken from urban, suburban, and natural scenes, covering diverse visual changes. Pitts30k Torii et al. (2013) extracted from Google Street View, mainly presents severe variations in viewpoint. Tokyo24/7 Torii et al. (2015) shows dramatic condition (light) changes. Nordland Sünderhauf et al. (2013) is gathered across four seasons with a fixed perspective from the front of a train. Moreover, we also use the Baidu Mall Sun et al. (2017), SPED Chen et al. (2017a), Pitts250k Torii et al. (2013), St. Lucia Glover et al. (2010), and SVOX Berton et al. (2021b) datasets for a few supplementary experiments.
@@
Performance Evaluation. We follow the previous work Berton et al. (2022a, b) using the Recall@N (R@N) as the evaluation metric for recognition performance. R@N is the proportion of queries for which at least one of the top-N predicted images is within a threshold of ground truth. We set the threshold to 10 frames for Nordland and 25 meters for others, as in this benchmark Berton et al. (2022b).